OpenStudy (anonymous):

You are the proud new owner of an Acme Deluxe Bread Kneading Machine, which you are using for the first time today. Suppose that at noon the dimensions of your loaf of bread dough are L=7 in, W=5 in, and H=4 in. At that time, you place the loaf in the machine for kneading and the machine begins by stretching the loaf's length at an initial rate of .75 in/min, punching down the loaf's height at a rate of 1 in/min, and increasing the loaf's width at a rate of .5 in/min. What is the rate of change of the volume of the loaf when the machine starts? Is the dough increasing or decreasing in size at

OpenStudy (anonymous):

\[V(t) = LWH = L(t)W(t)H(t)\] \[\frac{ dV }{ dt } = L'(t)(WH) + L \left( W'(t)H + H'(t)W \right)\]

OpenStudy (anonymous):

lmk if you have questions

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."